OpenStudy (anonymous):

y^5 * y^14 = a) 2y^19 b) y^60 c) y^19 d) 2y^60

OpenStudy (anonymous):

\(n^p*n^q=n^{p+q}\)

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

5^5 over 5^2 =

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you multiply, you add the exponents, but when you divide, you subtract exponents

OpenStudy (anonymous):

Oh ok Thank you so much you saved me today!

OpenStudy (anonymous):

you can actually see it quite easily if you write it out \[\frac{ 5*5*5*5*5 }{ 5*5 }\]

OpenStudy (anonymous):

all you will have left is 5*5*5